# Adaptive Collaborative Correlation Learning-based Semi-Supervised Multi-Label Feature Selection

### v3

@@ Abstract
Index Terms:
feature selection, semi-supervised multi-label learning, generalized regression model, adaptive similarity graph learning.
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ 1 Introduction
Multi-label data, where each instance is associated with multiple labels, widely exists in real-world applications [1, 2]. For example, a short video can be classified into several categorizes in the task of video classification [3], a news article may be related to several subjects in text categorization [4], and a painting image is possible to be tagged into a number of genres in the image annotation task [5]. In theses applications, multi-label data is often represented in a high-dimensional feature space with redundant and noisy features, resulting in the ”curse of dimensionality” problem and impacting the performance of subsequent tasks. Consequently, developing an effective method to reduce the dimensionality of multi-label data and thereby enhance downstream task performance has emerged as a pressing issue in practical applications.
@@
Multi-label feature selection (MFS) deals with this issue by choosing a compact subset of discriminative features from the original high-dimensional feature space, thus eliminating redundant and noisy information [6, 7]. In recent years, numerous MFS methods have been developed, which can be broadly categorized into two groups. The first group converts multi-label data into multiple independent single-label data and then applies the traditional feature selection method to the decomposed data [8, 9, 10]. Typical feature selection methods for single-label data include the Laplacian Score [9], spectral feature selection [11] and group lasso-based feature selection [12]. Those methods usually assess the importance of features based on specific criteria while maintaining the local manifold structure of data, and subsequently select the top ranked features. However, those methods fail to consider the correlations among different labels, which could enhance the performance of feature selection. Instead of decomposing multi-label data into several single-label data, the second group of MFS methods directly construct a model from multi-label data to facilitate feature selection [13, 14, 15, 16]. MIFS (multi-label informed feature selection) is a typical method in the second group, which incorporates multi-label information into a low-dimensional subspace to identify discriminative features [13]. Besides, Zhang et al. proposed using both the local label correlation and global label correlation to guide the feature selection process [17]. Moreover, Fan et al. have developed a collaborative learning framework for multi-label feature selection, which learns the label correlation and the feature correlation simultaneously [16]. Despite the demonstrated efficacy of these proposed methods in feature selection, they share a common underlying assumption, i.e., all data has been comprehensively labeled. Unfortunately, obtaining large amounts of labeled data is expensive and time-consuming in real-world applications [18]. A more common scenario is that only a portion of the samples are labeled [19, 20]. The aforementioned method based on a supervised way cannot be directly applied in the semi-supervised scenario. To address this issue, Chang et al. proposed a unified semi-supervised multi-label feature selection framework through the integration of sparse feature learning with label prediction [21]. Guo et al. developed a multi-label propagation mechanism to learn unknown labels, which was then integrated into an extended linear discriminant analysis for feature selection [22]. Lv et al. combined adaptive graph structure learning and manifold learning to investigate the global structure of data and label correlations, and subsequently embedded them into the feature selection process [23]. Constructing or learning the similarity graph of samples is a promising technology in MFS [24]. However, current semi-supervised multi-label feature selection methods either fail to investigate the sample similarity graph or rely exclusively on predefined graph approaches. Noise and outliers in the original feature space can undermine the reliability of predefined sample similarity graphs. Additionally, these methods describe label correlations in a pre-computed manner, which can lead to inaccuracies and deficiencies, especially with large volumes of unlabeled data. Existing methods that consider only the correlations between labels or samples overlook the potential for mutual enhancement, which can improve feature selection performance. Furthermore, these methods focus solely on the discriminative capability of selected features while ignoring their redundancy, ultimately degrading the performance of feature selection.
@@
To address these problems, we present a novel MFS method for semi-supervised multi-label data, named as Adaptive Collaborative Correlation lEarning-based Semi-Supervised Multi-label Feature Selection (Access-MFS). Specifically, the proposed method Access-MFS first embeds the feature selection process into a generalized regression model equipped with an extended uncorrelated constraint, which enables the selection of discriminative yet irrelevant features while preserving consistency between predicted and ground-truth labels in labeled data. We then integrate the instance correlation and label correlation into the proposed regression model to adaptively learn the sample similarity graph and label similarity graph. The adaptive collaborative correlation learning module is capable of preserving the sample similarity from high-dimensional to low-dimensional space, guaranteeing that similar samples are assigned similar labels and that strong label correlations lead to consistent predicted labels, simultaneously. It yields more accurate and reliable sample and label similarity graphs than those obtained through the predefined graph method. Finally, an iterative optimization algorithm is developed to solve our model. Extensive experimental results demonstrate that the proposed Access-MFS is superior to the state-of-the-art methods.
@@
To sum up, the main contributions of this paper are as follows: (i) To the best of our knowledge, we are the first to introduce adaptive collaborative correlation learning for feature selection on semi-supervised multi-label data. Our method can adaptively learn sample and label similarity graphs directly from the data, rather than relying on predefined graphs. The simultaneous learning of these two similarity graphs is integrated into the feature selection process, where they mutually enhance each other, leading to an improvement in feature selection performance. (ii) We propose a novel model for semi-supervised multi-label feature selection, incorporating a generalized regression module with an extended uncorrelated constraint and an adaptive collaborative correlation learning module. This model effectively selects features that are both discriminative and uncorrelated, while ensuring consistency between predicted labels and existing label data. (iii) We develop an efficient alternative optimization algorithm with guaranteed convergence to solve the proposed model. Comprehensive experiments are conducted on the real-world datasets to verify its superiority over state-of-the-art methods.
@@ 2 Related Work
Unlike the methods in the first category, the second category explores label correlations and constructs the feature selection model directly on multi-label data. MIFS selects discriminative features by factorizing the label matrix to reveal label correlations and diminish the impact of imperfect label information [13]. In GLFS [17], the LASSO-based sparse learning is used to select label-specific features that preserve groups, enabling the simultaneous exploration of label-group and instance-group correlations. However, these methods assume that all samples are labeled, which may not hold true in real-world scenarios. More often, only a subset of samples are labeled. Multi-label feature selection within the semi-supervised learning paradigm is proposed to address this issue. In LEDA [22], Guo et al. devised a semi-supervised multi-label propagation method to capture information from both labeled and unlabeled data simultaneously. Subsequently, the acquired label information is incorporated into an extended linear discriminant analysis module to select informative features. To explore the local manifold structures of both feature and label spaces in semi-supervised multi-label data, SMLFS [27] uses a local logistic regression model that incorporates feature graph regularization and label graph regularization. In addition, ${\ell_{2,p}}$-norm is applied to the regression coefficient matrix to select important feature dimensions. In [28], Li et al. proposed a semi-supervised multi-label feature selection method that simultaneously utilizes label and sample correlations. By imposing binary hash constraints on the spectral embedding model, SFS-BLL [29] identifies pseudo-labels for unlabeled data and utilizes a self-weighted sparse regression model to select discriminative features, leveraging both existing labels and learned pseudo-labels.
@@
However, the aforementioned semi-supervised multi-label feature selection methods are confronted with three issues. First, they use a predefined way to describe the sample similarity graph, which is easily influenced by noise and outliers. Second, these methods, which pre-compute label correlations based on partially known labels, fail to accurately capture the relationships between labels due to the large number of labels that remain unknown in a semi-supervised scenario. Existing methods fail to simultaneously consider the correlations between samples and between labels, overlooking how these two correlations can mutually enhance feature selection performance. Last, they only take into account the discriminative capability of the chosen features and overlook their redundancy, which will limit the performance of feature selection.
@@ 3.1 Notations
Let $\mathbf{X}=\left[\mathbf{X}_{l},\mathbf{X}_{u}\right]\in\mathbb{R}^{d\times n}$ be a given data matrix with $d$ dimensional features and $n$ instances, where $\mathbf{X}_{l}\in\mathbb{R}^{d\times n_{1}}$ is the labeled data, $\mathbf{X}_{u}\in\mathbb{R}^{d\times n_{2}}$ is the unlabeled data and ${n_{1}+n_{2}=n}$. The corresponding label matrix of $\mathbf{X}$ is denoted by $\mathbf{Y}=\left[\begin{array}[]{c}\mathbf{Y}_{l}\\ \mathbf{Y}_{u}\end{array}\right]\in\mathbb{R}^{n\times c}$ with $c$ categories, where $\mathbf{Y}_{l}\in\mathbb{R}^{n_{1}\times c}$ and $\mathbf{Y}_{u}\in\mathbb{R}^{n_{2}\times c}$ denote the labels of $\mathbf{X}_{l}$ and $\mathbf{X}_{u}$, respectively. In $\mathbf{Y}_{l}$, ${y_{ij}=1}$ indicates the $i$th instance is labeled as the $j$th class, otherwise it is 0. The label matrix $\mathbf{Y}_{u}$ of the unlabeled data is set to 00. In semi-supervised multi-label feature selection, our aim is to select $k$ informative features from the given dataset, which is much less than the total number of features in it. The framework of Access-MFS is shown in Fig. 1.
@@ 3.2 Formulation of Access-MFS
To select the discriminative and uncorrelated features, we propose an extended uncorrelated constraint on $\mathbf{W}$, denoted as $\mathbf{W}^{T}\mathbf{R}\mathbf{W}=\mathbf{I}$. Here, $\mathbf{R}$ is defined as $\mathbf{X}\mathbf{H}\mathbf{X}^{T}+\lambda\mathbf{D}+\theta\mathbf{X}\mathbf{L}_{s}\mathbf{X}^{T}$, where the matrix $\mathbf{D}\in\mathbb{R}^{d\times d}$ is diagonal, with the $i$th diagonal entry given by $\mathbf{D}(i,i)=\frac{1}{2\sqrt{\|\mathbf{w}_{i.}\|_{2}^{2}+\epsilon}}$ ($\epsilon$ is a small constant to avoid division by 0). Additional, $\mathbf{L}_{s}=\mathbf{D}_{s}-\mathbf{S}$ represents the Laplacian matrix of instance similarity matrix $\mathbf{S}$, $\mathbf{D}_{s}$ is the degree matrix of $\mathbf{S}$, and $\theta$ is a trade-off parameter. The uncorrelated constraint we proposed consists of three terms obtained by expanding $\mathbf{R}$. The first term aims to promote the orthonormality of the projected dimensions, which makes the low-dimensional instances to be uncorrelated. The second term is added to the first term to avoid the singularity of $\mathbf{X}\mathbf{H}\mathbf{X}^{T}$. The third term is advantageous in reducing the variance among samples within the same neighborhood under the graph structure. When $\lambda=\theta=0$, the proposed constraint reduces to the traditional uncorrelated constraint. The proposed constraint is beneficial to select discriminative features that are also uncorrelated.
@@
For semi-supervised multi-label feature selection [22, 23, 16, 28], the sample similarity structure and the label correlations serve as two crucial priors. Most existing methods use a predefined graph approach to capture the sample similarity or the label correlations. However, the presence of noise and outliers within the original feature space can undermine the reliability of the resulting graph. Besides, in semi-supervised settings where some labels are unknown, this approach cannot fully ascertain the label correlations. In this paper, we maintain the sample similarity structure and investigate label correlations by adaptively learning both the instance similarity graph and the label similarity graph. Guided by two fundamental assumptions that similar instances tend to have similar labels, and that samples close to each other in high-dimensional space should also be close in low-dimensional space, we formulate the instance similarity graph learning as follows:
$\min_{\mathbf{S}}\frac{1}{2}\sum_{i,j=1}^{n}\|\mathbf{f}_{i.}-\mathbf{f}_{j.}\|_{2}^{2}{s_{ij}}+\frac{1}{2}\sum_{i,j=1}^{n}\|\mathbf{W}^{T}\mathbf{x}_{.i}-\mathbf{W}^{T}\mathbf{x}_{.j}\|_{2}^{2}s_{ij}$ (3)
$~{}~{}~{}~{}~{}+\alpha\|\mathbf{S}\|_{F}^{2}$
${s.t.}~{}s_{ii}=0,s_{ij}\geq 0,\mathbf{S}\mathbf{1}_{n}=\mathbf{1}_{n},$
where $\mathbf{S}$ is the instance similarity matrix, $\mathbf{f}_{i.}$ in the predicted label matrix $\mathbf{F}$ denotes the label of the $i$th instance $\mathbf{x}_{.i}$, and $\alpha$ is a regularization parameter to control the sparsity of $\mathbf{S}$. The first two terms of Eq. (LABEL:New3) are designed to learn the instance similarity matrix from the perspectives of the label space and the feature space respectively. The third term serves as a regularization to prevent the trivial solution where each row in $\mathbf{S}$ contains only one element with a value of 1, while all other elements are 0. By defining the Laplacian matrix $\mathbf{L}_{s}=\mathbf{D}_{s}-\mathbf{S}$, where $\mathbf{D}_{s}$ is a diagonal matrix with its $i$th diagonal element equals to the sum of the $i$th row of $\mathbf{S}$, we can simplify Eq. (LABEL:New3) as follows:
$\min_{\mathbf{S}}\operatorname{Tr}(\mathbf{F}^{T}\mathbf{L}_{s}\mathbf{F})+\operatorname{Tr}(\mathbf{W}^{T}\mathbf{X}\mathbf{L}_{s}\mathbf{X}^{T}\mathbf{W})+\alpha\|\mathbf{S}\|_{F}^{2}$ (4)
${s.t.}~{}s_{ii}=0,s_{ij}\geq 0,\mathbf{S}\mathbf{1}_{n}=\mathbf{1}_{n}.$
Furthermore, in order to explore label correlations, we operate under the assumption that a stronger positive correlation between two labels leads to greater similarity in the predicted labels. This assumption guides the learning of the label similarity graph, which can be expressed as follows:
$\min_{\mathbf{P}}\frac{1}{2}\sum_{i,j=1}^{c}\|\mathbf{f}_{.i}-\mathbf{f}_{.j}\|_{2}^{2}p_{ij}+\beta\|\mathbf{P}\|_{F}^{2}$ (5)
${s.t.}~{}p_{ii}=0,p_{ij}\geq 0,\mathbf{P}\mathbf{1}_{c}=\mathbf{1}_{c},$
where $\mathbf{P}\in\mathbb{R}^{c\times c}$ denotes the label similarity matrix, $\mathbf{f}_{.i}$ is the $i$th label vector corresponding to the $i$th column of $\mathbf{F}$, and $\beta$ is a regularization parameter. From Eq. (5), it is evident that a larger $p_{ij}$ value indicates greater similarity between $\mathbf{f}_{.i}$ and $\mathbf{f}_{.j}$, and the converse is also true. We define the Laplacian matrix $\mathbf{L}_{p}=\mathbf{D}_{p}-\mathbf{P}$, where $\mathbf{D}_{p}$ is a diagonal matrix with the $i$th diagonal entry is $\sum_{j=1}^{c}p_{ij}$. Then, Eq. (5) can be rewritten as
${s.t.}~{}p_{ii}=0,p_{ij}\geq 0,\mathbf{P}\mathbf{1}_{c}=\mathbf{1}_{c}.$
By jointly utilizing Eqs. (LABEL:New4) and (6), we are able to collaboratively learn both the sample similarity matrix and the label similarity matrix in an adaptive manner, rather than relying on the predefined method. This is helpful to enhance the performance of feature selection.
@@
By combining Eqs. (2), (LABEL:New4) and (6) together, the proposed semi-supervised multi-label feature selection method Access-MFS is summarized as follows:
$\quad\quad+\theta(\operatorname{Tr}(\mathbf{F}^{T}\mathbf{L}_{s}\mathbf{F})+\operatorname{Tr}(\mathbf{W}^{T}\mathbf{X}\mathbf{L}_{s}\mathbf{X}^{T}\mathbf{W})+\alpha\|\mathbf{S}\|_{F}^{2})$
@@
where $\Xi=\{\mathbf{W},\mathbf{b},\mathbf{F},\mathbf{S},\mathbf{P}\}$, and $\theta$ and $\mu$ are two trade-off hyper-parameters. These two regularization parameters $\alpha$ and $\beta$ can be automatically determined during the optimization process.
In our Access-MFS method, we integrate feature selection, the collaborative learning of the instance similarity graph and the label similarity graph, and the label learning into a unified framework, enabling these different learning tasks to promote each other. Access-MFS can offer two key benefits: First, the proposed extended regression model not only maintains consistency between the predicted labels and the ground-truth of the labeled data but also selects discriminative and uncorrelated features. Second, Access-MFS is capable of adaptively learning the instance and label similarity graphs that are more reliable than those generated by predefined methods. This collaborative learning of the two similarity-induced graphs can preserve the local structures of the feature space and label space concurrently.

### v4

@@ Abstract
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
@@ 1 Introduction
Multi-labelMulti-label data, where each instance is associated with multiple labels, widely exists in real-world applications [1, 2]. For example, a short video can be classified into several categorizes in the task of video classification [3], a news article may be related to several subjects in text categorization [4], and a painting image is possible to be tagged into a number of genres in the image annotation task [5]. In these applications, multi-label data is often represented in a high-dimensional feature space with redundant and noisy features, resulting in the ”curse of dimensionality” problem and impacting the performance of subsequent tasks. Consequently, developing an effective method to reduce the dimensionality of multi-label data and thereby enhance downstream task performance has emerged as a pressing issue in practical applications.
@@
Multi-label feature selection (MFS) deals with this issue by choosing a compact subset of discriminative features from the original high-dimensional feature space, thus eliminating redundant and noisy information [6, 7]. In recent years, numerous MFS methods have been developed, which can be broadly categorized into two groups. The first group converts multi-label data into multiple independent single-label data and then applies the traditional feature selection method to the decomposed data [8, 9, 10]. Typical feature selection methods for single-label data include the Laplacian Score [9], spectral feature selection [11] and group lasso-based feature selection [12]. Those methods usually assess the importance of features based on specific criteria while maintaining the local manifold structure of data, and subsequently select the top ranked features. However, those methods fail to consider the correlations among different labels, which could enhance the performance of feature selection. Instead of decomposing multi-label data into several single-label data, the second group of MFS methods directly construct a model from multi-label data to facilitate feature selection [13, 14, 15, 16]. MIFS [13] is a typical method in the second group, which incorporates multi-label information into a low-dimensional subspace to identify discriminative features. Besides, Zhang et al. proposed using both local label correlations and global label correlations to guide the feature selection process [17]. Moreover, Fan et al. have developed a collaborative learning framework for multi-label feature selection, which learns the label correlations and the feature correlations simultaneously [16]. Despite the demonstrated efficacy of these proposed methods in feature selection, they share a common underlying assumption, i.e., all data has been comprehensively labeled. Unfortunately, obtaining large amounts of labeled data is expensive and time-consuming in real-world applications [18]. A more common scenario is that only a portion of the samples are labeled [19, 20]. The aforementioned method based on a supervised way cannot be directly applied in the semi-supervised scenario. To address this issue, Chang et al. proposed a unified semi-supervised multi-label feature selection framework through the integration of sparse feature learning with label prediction [21]. Guo et al. developed a multi-label propagation mechanism to learn unknown labels, which was then integrated into an extended linear discriminant analysis for feature selection [22]. Lv et al. combined adaptive graph structure learning and manifold learning to investigate the global structure of data and label correlations, and subsequently embedded them into the feature selection process [23]. Constructing or learning the similarity graph is a promising technology for enhancing learning performance in MFS [24]. However, existing semi-supervised MFS methods rely on predefined sample similarity graphs, while noise and outliers in the original feature space can undermine their reliability. Moreover, the label correlations encoded in a pre-computed manner may introduce inaccuracies and deficiencies, especially with large volumes of unlabeled data. Additionally, current methods that consider either label or sample correlations often fail to exploit their mutual reinforcement, thereby limiting the potential for enhanced feature selection. Furthermore, these methods focus solely on the discriminative capability of selected features while ignoring their redundancy, ultimately degrading the performance of feature selection.
@@
To address these limitations, we present a novel MFS method for semi-supervised multi-label data, named as Adaptive Collaborative Correlation lEarning-based Semi-Supervised Multi-label Feature Selection (Access-MFS). Specifically, the proposed method Access-MFS first embeds the feature selection process into a generalized regression model equipped with an extended uncorrelated constraint, which enables the selection of discriminative yet irrelevant features while preserving consistency between predicted and ground-truth labels in labeled data. We then integrate the instance correlations and label correlations into the proposed regression model to adaptively learn the sample similarity graph and label similarity graph. The adaptive collaborative correlation learning module is capable of preserving the sample similarity from high-dimensional to low-dimensional space, guaranteeing that similar samples are assigned similar labels and that strong label correlations lead to consistent predicted labels, simultaneously. It yields more accurate and reliable sample and label similarity graphs than those obtained through the predefined graph method. Finally, an iterative optimization algorithm is developed to solve our model. Extensive experimental results demonstrate that the proposed Access-MFS is superior to the state-of-the-art methods.
@@
To sum up, the main contributions of this paper are as follows: (i) To the best of our knowledge, we are the first to introduce adaptive collaborative correlation learning for feature selection on semi-supervised multi-label data. Our method can adaptively learn sample and label similarity graphs directly from the data, rather than relying on predefined graphs. The simultaneous learning of these two similarity graphs is integrated into the feature selection process, where they mutually enhance each other, leading to improved feature selection performance. (ii) We propose a novel model for semi-supervised multi-label feature selection, incorporating a generalized regression module with an extended uncorrelated constraint and an adaptive collaborative correlation learning module. This model effectively selects features that are both discriminative and uncorrelated, while ensuring consistency between predicted labels and existing label data. (iii) We develop an efficient alternative optimization algorithm with guaranteed convergence to solve the proposed model. Comprehensive experiments are conducted on the real-world datasets to verify its superiority over state-of-the-art methods.
@@ 2 Related Work
Unlike the methods in the first category, the second category explores label correlations and constructs the feature selection model directly on multi-label data. MIFS selects discriminative features by factorizing the label matrix to reveal label correlations and diminish the impact of imperfect label information [13]. In GLFS [17], the LASSO-based sparse learning is used to choose label-specific features that preserve groups, allowing for the simultaneous exploration of label-group and instance-group correlations. However, these methods assume that all samples are labeled, which may not hold true in real-world scenarios. More often, only a subset of samples are labeled. Multi-label feature selection within the semi-supervised learning paradigm is proposed to address this issue. For instance, Guo et al. devised a semi-supervised multi-label propagation method to capture information from both labeled and unlabeled data simultaneously. Subsequently, the acquired label information is incorporated into an extended linear discriminant analysis module to select informative features [22]. Xu et al. developed a semi-supervised multi-label feature selection algorithm that utilizes the probabilistic neighborhood similarities to learn the correlations between features and maintains consistent correlation information within the label space [27]. To explore the local manifold structures of both feature and label spaces in semi-supervised multi-label data, Xu et al. [28] uses a local logistic regression model that incorporates feature graph regularization and label graph regularization. In addition, $\ell_{2,p}$}-norm is applied to the regression coefficient matrix to select important feature dimensions. Lv et al. [23] introduced two modules to accurately represent the global structure of data and its label correlations: the adaptive global structure learning module, which maintains a global and sparse reconstruction of the structure, and the manifold learning module, focused on capturing the local structure and label correlations. These modules work synergistically to address their respective limitations, facilitating the selection of discriminative features. In SMDR-IC [29], Li et al. proposed a semi-supervised multi-label feature selection method that simultaneously utilizes label and sample correlations. By imposing binary hash constraints on the spectral embedding model, the method named SFS-BLL [30] has been proposed which identifies pseudo-labels for unlabeled data and utilizes a self-weighted sparse regression model to select discriminative features, leveraging both existing labels and learned pseudo-labels. Furthermore, Li et al. extended SMDR-IC and proposed SMDR-MRC [31], which decomposes the feature selection matrix into two matrices to respectively identify specific and common features, and introduces a non-zero correlation constraint to reduce redundant information between them. Using a shared subspace learning paradigm, Sheikhpour et al. [32] decomposed the feature selection matrix into a label-specific weight matrix and a shared feature subspace term, while preserving both label relationships and feature correlations.Qing et al. [33] proposed a semi-supervised multi-label feature selection method that selects label-relevant and non-redundant features using clustering, scoring, and ranking aggregation.
@@
Although the aforementioned semi-supervised multi-label feature selection methodshave demonstrates certain effectiveness, several challenges still remain unaddressed. First, they use a predefined way to describe the sample similarity graph, which is easily influenced by noise and outliers. Second, these methods, which pre-compute label correlations based on partially known labels, fail to accurately capture the relationships between labels due to the large number of labels that remain unknown in a semi-supervised scenario. And existing methods fail to simultaneously consider the correlations between samples and between labels, overlooking how these two correlations can mutually enhance feature selection performance. Last, most approaches only take into account the discriminative capability of the chosen features and overlook their redundancy, which will limit the performance of feature selection. These issues are exhuastively discussed in our proposed Access-MFS method.
@@ 3.1 Notations
Let $\mathbf{X}=\left[\mathbf{X}_{l},\mathbf{X}_{u}\right]\in\mathbb{R}^{d\times n}$ be a given data matrix with $d$ dimensional features and $n$ instances, where $\mathbf{X}_{l}\in\mathbb{R}^{d\times n_{1}}$ is the labeled data, $\mathbf{X}_{u}\in\mathbb{R}^{d\times n_{2}}$ is the unlabeled data and $n_{1}+n_{2}=n$}. The corresponding label matrix of $\mathbf{X}$ is denoted by $\mathbf{Y}=\left[\begin{array}[]{c}\mathbf{Y}_{l}\\ \mathbf{Y}_{u}\end{array}\right]\in\mathbb{R}^{n\times c}$ with $c$ categories, where $\mathbf{Y}_{l}\in\mathbb{R}^{n_{1}\times c}$ and $\mathbf{Y}_{u}\in\mathbb{R}^{n_{2}\times c}$ denote the labels of $\mathbf{X}_{l}$ and $\mathbf{X}_{u}$, respectively. In $\mathbf{Y}_{l}$, $y_{ij}=1$} indicates the $i$th instance is labeled as the $j$th class, otherwise it is 0. The label matrix $\mathbf{Y}_{u}$ of the unlabeled data is set to 0. In semi-supervised multi-label feature selection, our aim is to select $k$ informative features from the given dataset, which is much less than the total number of features in it. The framework of Access-MFS is shown in Fig. 1.
@@ 3.2 Formulation of Access-MFS
To select the discriminative and uncorrelated features, we propose an extended uncorrelated constraint on $\mathbf{W}$, denoted as $\mathbf{W}^{T}\mathbf{R}\mathbf{W}=\mathbf{I}$. Here, $\mathbf{R}$ is defined as $\mathbf{X}\mathbf{H}\mathbf{X}^{T}+\lambda\mathbf{D}+\theta\mathbf{X}\mathbf{L}_{s}\mathbf{X}^{T}$, where the matrix $\mathbf{D}\in\mathbb{R}^{d\times d}$ is diagonal, with the $i$th diagonal entry given by $\mathbf{D}(i,i)=\frac{1}{2\sqrt{\|\mathbf{w}_{i.}\|_{2}^{2}+\epsilon}}$ ($\epsilon$ is a small constant to avoid division by 0). Additional, $\mathbf{L}_{s}=\mathbf{D}_{s}-\mathbf{S}$ represents the Laplacian matrix of instance similarity matrix $\mathbf{S}$, $\mathbf{D}_{s}$ is the degree matrix of $\mathbf{S}$, and $\theta$ is a trade-off parameter. The uncorrelated constraint we proposed consists of three terms obtained by expanding $\mathbf{R}$. The first term aims to promote the orthonormality of the projected dimensions, which makes the low-dimensional instances to be uncorrelated. The second term is added to the first term to avoid the singularity of $\mathbf{X}\mathbf{H}\mathbf{X}^{T}$. The third term is advantageous in reducing the variance among samples within the same neighborhood under the graph structure. When $\lambda=\theta=0$, the proposed constraint reduces to the traditional uncorrelated constraint. The proposed constraint is beneficial to select discriminative features that are also uncorrelated. Furthermore, it can simplify the optimization process of $\mathbf{W}$ and transforms $\mathbf{W}$ into a closed-form solution, as demonstrated in the optimization procedure.
@@
The exploration of sample similarity structure and label correlations plays an important role in enhancing the performance of semi-supervised multi-label feature selection [16, 22, 23, 29]. Many existing methods capture sample similarity and label correlations using predefined graph approaches. However, the presence of noise and outliers in the original feature space can compromise the accuracy of the resulting graph. Moreover, in semi-supervised settings with partially unknown labels, this approach cannot fully ascertain label correlations. In this paper, we preserve the sample similarity structure and capture label correlations by adaptively learning both the instance similarity graph and the label similarity graph. This objective can be formally expressed as follows:
$\min_{\mathbf{S},\mathbf{P}}\operatorname{Tr}(\mathbf{W}^{T}\mathbf{X}\mathbf{L}_{s}\mathbf{X}^{T}\mathbf{W})+\operatorname{Tr}(\mathbf{F}^{T}\mathbf{L}_{s}\mathbf{F})+\alpha\|\mathbf{S}\|_{F}^{2}$ (3)
$+\operatorname{Tr}(\mathbf{F}\mathbf{L}_{p}\mathbf{F}^{T})+\beta\|\mathbf{P}\|_{F}^{2}$
${s.t.}~s_{ii}=0,s_{ij}\geq 0,\mathbf{S}\mathbf{1}_{n}=\mathbf{1}_{n},p_{ii}=0,p_{ij}\geq 0,\mathbf{P}\mathbf{1}_{c}=\mathbf{1}_{c},$
where $\mathbf{S}$ and $\mathbf{P}$ denote the instance and label similarity matrix, respectively, $\mathbf{L}_{s}=\mathbf{D}_{s}-\mathbf{S}$ and $\mathbf{L}_{p}=\mathbf{D}_{p}-\mathbf{P}$ are their corresponding Laplacian matrices, with $\mathbf{D}_{s}$ and $\mathbf{D}_{p}$ being diagonal matrices whose diagonal entries are $\sum_{j=1}^{n}s_{ij}$ and $\sum_{j=1}^{c}p_{ij}$. Regularization parameters $\alpha$ and $\beta$ are used to control the sparsity of two similarity matrices. The first two terms of Eq. (LABEL:New3) are designed to learn the instance similarity matrix from the perspectives of the feature space and the label space, respectively, while the forth term helps to maintain strong label correlations by encouraging related labels to have similar predicted values. The Frobenius regularization terms $\alpha\|\mathbf{S}\|_{F}^{2}$} and $\beta\|\mathbf{P}\|_{F}^{2}$} prevent trivial solutions where each row of $\mathbf{S}$ or $\mathbf{P}$ collapses to a one-hot vector. By optimizing Eq. (LABEL:New3), our framework collaboratively and adaptively learns both the instance similarity graph and the label similarity graph, rather than relying on predefined structures. This bidirectional coupling enables the two graphs to iteratively refine and reinforce each other, thus making them less vulnerable to noise, outliers, and label sparsity. As a result, the learned graphs provide more reliable structural priors, which substantially enhance the performance of semi-supervised multi-label feature selection.
@@
By combining Eqs. (2) and (LABEL:New3) together, the proposed semi-supervised multi-label feature selection method Access-MFS is summarized as follows:
$\min_{\Xi}\|\mathbf{X}^{T}\mathbf{W}+\mathbf{1}_{n}\mathbf{b}^{T}-\mathbf{F}\|_{F}^{2}+\lambda\|\mathbf{W}\|_{2,1}+{\|{\mathbf{F}_{l}-\mathbf{Y}_{l}}\|_{F}^{2}}$ (4)
$\quad\quad+\theta(\operatorname{Tr}(\mathbf{W}^{T}\mathbf{X}\mathbf{L}_{s}\mathbf{X}^{T}\mathbf{W})+\operatorname{Tr}(\mathbf{F}^{T}\mathbf{L}_{s}\mathbf{F})+\alpha\|\mathbf{S}\|_{F}^{2})$
@@
where $\Xi=\{\mathbf{W},\mathbf{b},\mathbf{F},\mathbf{S},\mathbf{P}\}$, and $\theta$ and $\mu$ serve as trade-off hyper-parameters to control how strongly instance correlation and label correlation are captured during learning. These two regularization parameters $\alpha$ and $\beta$ can be automatically determined during the optimization process.
In our Access-MFS method, we integrate feature selection, the collaborative learning of instance similarity graphs and label similarity graphs, and the label learning into a unified framework, enabling these different learning tasks to promote each other. Access-MFS can offer two key benefits: First, the proposed extended regression model not only maintains consistency between the predicted labels and the ground-truth of the labeled data but also selects discriminative and uncorrelated features. Second, Access-MFS is capable of adaptively learning the instance and label similarity graphs that are more reliable than those generated by predefined methods. This collaborative learning of the two similarity-induced graphs can preserve the local structures of the feature space and label space concurrently.